# Can Knowledge Editing Really Correct Hallucinations?

### v2

@@ 5 Conclusion
In this paper, we have built a new benchmark HalluEditBench to holistically assess diverse knowledge editing techniques in correcting real-world hallucinations. First, we carefully construct a massive hallucination dataset based on Wikipedia with $9$ domains, $26$ topics, and more than $6,000$ hallucinations. Then, we systematically investigate the performance of different knowledge editing methods from five perspectives including Efficacy, Generalization, Portability, Locality, and Robustness. With extensive empirical experiment results, we have offered valuable insights for the potentials and limitations of typical knowledge editing methods, which could inspire more future improvements.
@@ Appendix A Reproducibility Statement
We conduct the experiments on eight NVIDIA RTX A6000 GPUs. The decoding temperatures are 00 to ensure the reproducibility. The model checkpoints are downloaded from https://huggingface.co/. The specific download links are as follows:
@@ Appendix B Details of the Benchmarked Knowledge Editing Techniques
FT-L (Meng et al., 2022) Constrained Fine-Tuning (FT-L) is a targeted approach to fine-tuning that focuses on adjusting a specific layer within a model’s feed-forward network (FFN). Guided by causal tracing results from ROME, FT-L modifies the layer most associated with the desired changes. The goal of FT-L is to fine-tune the model by maximizing the likelihood of the target sequence, particularly focusing on the prediction of the last token, ensuring that the model adapts to modified facts without affecting its broader performance. To achieve this, explicit parameter-space norm constraints are applied to the weights, ensuring minimal interference with unmodified facts and preserving the integrity of the model’s original knowledge.
FT-M (Zhang et al., 2024e) In contrast to FT-L, which fine-tunes by maximizing the probability of all tokens in the target sequence based on the last token’s prediction, Fine-Tuning with Masking (FT-M) refines this approach to align more closely with the traditional fine-tuning objective. FT-M also targets the same FFN layer identified by causal tracing but employs a masked training strategy. Specifically, it uses cross-entropy loss on the target answer while masking out the original text, ensuring that the model is trained directly on the relevant target content. This approach mitigates potential deviations from the original fine-tuning objective and provides a more precise adjustment of the model’s weights with minimal disruption to unrelated model behavior.
@@ Appendix C A More Detailed Related Work
Knowledge Editing has been adopted as one of the mainstream paradigms to address the hallucinations in LLMs efficiently (Chen & Shu, 2024a; Tonmoy et al., 2024). Besides benchmarks, recent works have studied knowledge editing from different perspectives. The first line of works aims to probe into the relationship between localization and editing and gain a deeper understanding of the working mechanisms of different techniques (Wang et al., 2024b; Niu et al., 2024; Hase et al., 2024a; b; Ferrando et al., 2024; Gupta et al., 2024; Chen et al., 2024d; c; Zou et al., 2023; Yao et al., 2024). For example, Hase et al. (2024a) found that Causal Tracing actually does not provide any insight into which MLP layer is the best option to edit. The second line of works intends to enhance the performance and applicability of knowledge editing in specific scenarios (Rozner et al., 2024; Jiang et al., 2024a; b; Zhang et al., 2024c; b; d; a; f; Wu et al., 2024; Qi et al., 2024; Sharma et al., 2024a; Li et al., 2024c; b; Fang et al., 2024; Wang & Li, 2024a; b; Wang et al., 2024g; f; d; 2023b; Cheng et al., 2024b; a; Xie et al., 2024; Bi et al., 2024c; b; a; Chen et al., 2024b; Wei et al., 2024b; Fei et al., 2024; Xu et al., 2024; Gu et al., 2023; Yin et al., 2024; Cai et al., 2024a; Liu et al., 2024b; Ge et al., 2024b; Deng et al., 2024; Peng et al., 2024). For example, Ma et al. (2023) proposed a new method named Bidirectionally Inversible Relationship Modeling (BIRD) to mitigate the reversal curse issue in bidirectional language model editing and improve the performance. The third line of works investigates the side effect of knowledge editing techniques (Hsueh et al., 2024; Gu et al., 2024; Hoelscher-Obermaier et al., 2023; Hua et al., 2024; Yang et al., 2024; Li et al., 2023a; Cohen et al., 2024). For example, Yang et al. (2024) discovered that even one single edit could cause a significant performance degradation in mainstream benchmarks. The fourth line of works explores the potential misuse risks of knowledge editing or its applications beyond correcting hallucinations (Chen et al., 2024a; Uppaal et al., 2024; Wang et al., 2024c; Cai et al., 2024b; Yan et al., 2024). For example, Chen et al. (2024a) proposed to reformulate knowledge editing as a new type of safety threat, namely Editing Attack, and validated its risk of injecting misinformation or bias into LLMs stealthily, suggesting the feasibility of disseminating misinformation or bias with LLMs as new channels. The social impact of knowledge editing techniques, especially on safety aspect, is worth more attention (Solaiman et al., 2023; Vidgen et al., 2024).
@@ Appendix D Impact Statement
Misinformation is a longstanding threat for online safety and public trust (Chen et al., 2022; Wang et al., 2023a). The conventional countermeasures include detection (Shu et al., 2017; Nan et al., 2024; 2023; Liu et al., 2024a), intervention (Bak-Coleman et al., 2022; Aghajari et al., 2023; Hartwig et al., 2024; Yue et al., 2024; He et al., 2023) and attribution (Huang et al., 2024a; b; Beigi et al., 2024). Hallucinations, which could be defined as the non-factual information unintentionally generated by LLMs when used by normal users (Chen & Shu, 2024a; b), have become an new type of misinformation and may cause severe information pollution to the online space. Knowledge editing is a promising paradigm to correct hallucinations and contribute to the fight against the misinformation crisis in the era of LLMs, due to its advantage of avoiding retraining from scratch. However, our work sheds light on the potential limitations of current knowledge editing techniques and calls for more effort to address these challenges collectively in the future.
@@ E.1 Generalization Scores of Knowledge Editing Methods on Each Domain
Figure 8: Generalization Scores of Knowledge Editing Methods on 3 LLMs and 2 Domains. Generalization Scores (%) are measured by the accuracy on five types of Generalization Evaluation Question-answer Pairs including Rephrased Questions (“rephrase”), two types of Yes-or-No Questions with Yes or No as answers (“yes” or “no”), Multi-Choice Questions (“mc”), Reversed Questions (“reversed”). The “average” refers to the averaged scores over five types of questions. The domains include “places” and “human”.

### v3

@@ 5 Conclusion
In this paper, we have built a new benchmark HalluEditBench to holistically assess diverse knowledge editing techniques in correcting real-world hallucinations. First, we meticulously construct a massive and comprehensive hallucination dataset based on Wikidata with $9$ domains, $26$ topics, and more than $6,000$ hallucinations. Then, we systematically investigate the performance of different knowledge editing methods from five perspectives including Efficacy, Generalization, Portability, Locality, and Robustness. Our findings reveal that previous benchmarks cannot reflect the true effectiveness of knowledge editing methods in correcting real-world hallucinations and current editing methods mostly show limited performance across five dimensions. We also offer valuable and actionable insights to inspire future advancements in knowledge editing for large language models.
Acknowledgments
This material is based upon work supported by the U.S. Department of Homeland Security under Grant Award Number 17STQAC00001-07-04, NSF awards (SaTC-2241068, IIS-2506643, and POSE-2346158), a Cisco Research Award, and a Microsoft Accelerate Foundation Models Research Award. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security and the National Science Foundation.
@@ Appendix A Reproducibility Statement
We conduct the experiments on NVIDIA RTX A6000 GPUs. The decoding temperatures are 00 to ensure the reproducibility. The model checkpoints are downloaded from https://huggingface.co/. The specific download links are as follows:
@@ Appendix B Details of the Benchmarked Knowledge Editing Techniques
FT-L (Zhu et al., 2020; Meng et al., 2022) Constrained Fine-Tuning (FT-L) is a targeted approach to fine-tuning that focuses on adjusting a specific layer within a model’s feed-forward network (FFN). Guided by causal tracing results from ROME, FT-L modifies the layer most associated with the desired changes. The goal of FT-L is to fine-tune the model by maximizing the likelihood of the target sequence, particularly focusing on the prediction of the last token, ensuring that the model adapts to modified facts without affecting its broader performance. To achieve this, explicit parameter-space norm constraints are applied to the weights, ensuring minimal interference with unmodified facts and preserving the integrity of the model’s original knowledge.
FT-M (Zhang et al., 2024f) In contrast to FT-L, which fine-tunes by maximizing the probability of all tokens in the target sequence based on the last token’s prediction, Fine-Tuning with Masking (FT-M) refines this approach to align more closely with the traditional fine-tuning objective. FT-M also targets the same FFN layer identified by causal tracing but employs a masked training strategy. Specifically, it uses cross-entropy loss on the target answer while masking out the original text, ensuring that the model is trained directly on the relevant target content. This approach mitigates potential deviations from the original fine-tuning objective and provides a more precise adjustment of the model’s weights with minimal disruption to unrelated model behavior.
@@ Appendix C A More Detailed Related Work
Knowledge Editing has been adopted as one of the mainstream paradigms to address the hallucinations in LLMs efficiently (Chen & Shu, 2024a; Tonmoy et al., 2024; Li et al., 2024e). Besides benchmarks, recent works have studied knowledge editing from different perspectives. The first line of works aims to probe into the relationship between localization and editing and gain a deeper understanding of the working mechanisms of different techniques (Wang et al., 2024b; Niu et al., 2024; Hase et al., 2024a; b; Ferrando et al., 2024; Gupta et al., 2024; Chen et al., 2024e; d; Zou et al., 2023; Yao et al., 2024; Wu et al., 2025). For example, Hase et al. (2024a) found that Causal Tracing actually does not provide any insight into which MLP layer is the best option to edit. The second line of works intends to enhance the performance and applicability of knowledge editing in specific scenarios (Rozner et al., 2024; Jiang et al., 2024a; b; Zhang et al., 2024d; c; e; b; g; 2025a; 2025b; Wu et al., 2024; Qi et al., 2024; Sharma et al., 2024a; Li et al., 2024c; b; Fang et al., 2024; Wang & Li, 2024a; b; Wang et al., 2024g; f; d; 2023b; Cheng et al., 2024b; a; Xie et al., 2024; Bi et al., 2024c; b; a; Chen et al., 2024c; b; Wei et al., 2024b; Fei et al., 2024; Xu et al., 2024; Gu et al., 2023; Yin et al., 2024; Cai et al., 2024a; Liu et al., 2024b; 2025; Ge et al., 2024b; Deng et al., 2024; Peng et al., 2024; Zhao et al., 2025; Jiang et al., 2025; Li et al., 2025; Lu et al., 2024; Zeng et al., 2024; Gu et al., 2025). For example, Ma et al. (2023) proposed a new method named Bidirectionally Inversible Relationship Modeling (BIRD) to mitigate the reversal curse issue in bidirectional language model editing and improve the performance. The third line of works investigates the side effect of knowledge editing techniques (Hsueh et al., 2024; Gu et al., 2024; Hoelscher-Obermaier et al., 2023; Hua et al., 2024; Yang et al., 2024a; b; Li et al., 2023a; Cohen et al., 2024). For example, Yang et al. (2024a) discovered that even one single edit could cause a significant performance degradation in mainstream benchmarks. The fourth line of works explores the potential misuse risks of knowledge editing or its applications beyond correcting hallucinations (Chen et al., 2024a; Uppaal et al., 2024; Wang et al., 2024c; Cai et al., 2024b; Yan et al., 2024; Zhang et al., 2025c; Grimes et al., 2024; Li et al., 2024d; Youssef et al., 2024; 2025). For example, Chen et al. (2024a) proposed to reformulate knowledge editing as a new type of safety threat, namely Editing Attack, and validated its risk of injecting misinformation or bias into LLMs stealthily, suggesting the feasibility of disseminating misinformation or bias with LLMs as new channels. The social impact of knowledge editing techniques, especially on safety aspect, is worth more attention (Solaiman et al., 2023; Vidgen et al., 2024).
@@ Appendix D Impact Statement
Misinformation is a longstanding threat for online safety and public trust (Chen et al., 2022; Wang et al., 2023a). The conventional countermeasures include detection (Shu et al., 2017; Nan et al., 2024; 2023; Liu et al., 2024a), intervention (Bak-Coleman et al., 2022; Aghajari et al., 2023; Hartwig et al., 2024; Yue et al., 2024; He et al., 2023) and attribution (Huang et al., 2024a; b; Beigi et al., 2024). Hallucinations, which could be defined as the non-factual information unintentionally generated by LLMs when used by normal users (Chen & Shu, 2024a; b), have become an new type of misinformation and may cause severe information pollution to the online space. Besides methods such as Retrieval-Augmented Generation (Shi et al., 2025; Ni et al., 2025; Zhou et al., 2024), knowledge editing is a promising paradigm to correct hallucinations and contribute to the fight against the misinformation crisis in the era of LLMs, due to its advantage of avoiding retraining from scratch. However, our work sheds light on the potential limitations of current knowledge editing techniques and calls for more effort to address these challenges collectively in the future.
@@ E.1 Generalization Scores of Knowledge Editing Methods on Each Domain
Figure 8: Generalization Scores of Knowledge Editing Methods on 3 LLMs and 2 Domains. Generalization Scores (%) are measured by the accuracy on five types of Generalization Evaluation Question-answer Pairs including Rephrased Questions (“rephrase”), two types of Yes-or-No Questions with Yes or No as answers (“yes” or “no”), Multi-Choice Questions (“mc”), Reversed Questions (“reversed”). The “average” refers to the averaged scores over five types of questions. The domains include “human” and “places”.